OpenStudy (anonymous):

Solve 2(3x – 4) = 3x + 7. You must show all work.

OpenStudy (anonymous):

Expand brackets and get 6x -8= 3x +7 Bring 3x to the left and -8 to the right This will leave you with 6x-3x=7 +8 simplifying gives you 3x=15 divide out 3 and you have that x=5 This is right because if you sub 5 for x in both equations, you get 2(15-4)= 2(11)=22 and 3(5) +7, which also equals 22

OpenStudy (anonymous):

ty thats helps alot

OpenStudy (anonymous):

my final question is this

OpenStudy (anonymous):

Solve K = 1/3 (b – 2a) for b. You must show all work.

OpenStudy (anonymous):

Im assuming you dont have some sort of b= random number or anything like that So you'd just do: K =1/3 (b – 2a) 3K = (b – 2a) (multiply both sides by 3 to get rid of the fraction) 3K + 2a = b (drop the brackets and bring the negative 2a to the right) b = 3K + 2a (B is therefore equal to whatever 3k is equal to + 2a)

OpenStudy (anonymous):

Sorry to the left for the second part of the answer*

OpenStudy (anonymous):

Part 1: Show the work to solve |x+3| +7 10 (3 points)

OpenStudy (anonymous):

does that mean that (x+3) +7 = 10? If that's what it is, there can two answers for the question, so I'm assuming the second part is some sort of "Describe x of the graph" or something So you can basically get: x+3+7=10 x+10=10 x=0 AND -x-3+7=10 -x+4=10 -x=6 x=-6 Because [x+3] can equal both x +3, or (-x+3) If theres a graphing part involved with the second part, you can say that X is a vertical line at the origin (0) OR its a vertical line at -6 (x-axis)